OpenStudy (anonymous):

convert decimal to octal for : 1. 629 2.351 win cendol and tempe and the bakso i treats

OpenStudy (anonymous):

teach me please

OpenStudy (anonymous):

well, first do you know how, binary, octal, base-3 so on work?

OpenStudy (saifoo.khan):

1629 = 3135

OpenStudy (anonymous):

1. 1165

OpenStudy (anonymous):

512*1+64*1+8*6+1*5=629

OpenStudy (anonymous):

629

OpenStudy (anonymous):

oh i mean : 625 and 351 sorry for de typo

OpenStudy (saifoo.khan):

LOOOOL.

OpenStudy (saifoo.khan):

629 >> 1165

OpenStudy (anonymous):

im serious now khan

OpenStudy (saifoo.khan):

351 >> 537

OpenStudy (anonymous):

537 5*64+3*8+7*1=351

OpenStudy (anonymous):

629 convert to octal 351 convert to octal

OpenStudy (saifoo.khan):

we already did.

OpenStudy (anonymous):

629=1165 (octal) 351=537 (octal)

OpenStudy (anonymous):

the detail steps

OpenStudy (anonymous):

629/8 = 78 sisa 5 78/8 = 9 sisa 6 9/8= 1 sisa 1 1/8 = 0 sisa 1 so the octal is 1165 thats the steps yah

OpenStudy (anonymous):

saifoo what your steps please

OpenStudy (anonymous):

emm i got difficult in english to describe

OpenStudy (anonymous):

first, lets break down a decimal number. lets look at the number 1203 1203= 1000+200+3 another way to write this (you will see the purpose soon) is: 1203= 1*(10^3)+2*(10^2)+0*(10^1)+3*(10^0) Now when you look at the same number in octal: we can break it down like this: 1203= 1024 + 128 + 48 + 3 now, to the exponent step: 1203= 2*(8^3)+2*(8^2)+6*(8^1)+3*(8^0) thus; 1203 (decimal) = 2263 (octal) Hopefully, now you see the idea of octal. So, now lets convert a number step by step! How about... 1035 (decimal) 8^4=4096 since 4096>1035 we know that the number in octal will only have 4 numbers in it. 8^3=512 1035/512=2.0214... so in the 8^3's place we will put a 2. 512*2=1024 1035-1024=11 since 64>11, the 8^2's place will have a zero. 11/8=1.375 so, there will be a 1 in the 8^1's place now we have 3 left! so there will be a 3 in the 8^0's spot so: 1053 (decimal)= 2013 (octal) hopefully that breakdown helped a little bit. different bases are hard to get a hold of because they are so counter-intuitive. Let me know if you have anymore questions.

OpenStudy (anonymous):

thank you so much shosny you bright my day, now i understand it

OpenStudy (anonymous):

Any time noras!